OpenStudy (anonymous):

A rectangular counter 3 feet wide and 5 feet long has a circle cut out of it in order to have a sink installed. The circle has a diameter of 18 inches. What is the approximate area of the remaining countertop surface? Use 3.14 for pi.

OpenStudy (anonymous):

3 feet = 36 inches 5 feet = 60 inches 2160 - (pi(9)^2) = 2160 - (3.14(81)) = 1905.66

OpenStudy (anonymous):

Would you say its 18.9. From any of these options which one would you choose 18.9,15.0,29.5,13.2

OpenStudy (anonymous):

\[3*5-\left(3.14*\left(\frac{18/12}{2}\right)^2\right)=13.2338 \]